OpenStudy (anonymous):

A rectangle is formed by joining two squares with area 49 cm2. What is the perimeter of the rectangle?

OpenStudy (asnaseer):

if the area of a square is 49 cm2 then what do you think its side length would be?

OpenStudy (anonymous):

my bad asnaseer

OpenStudy (asnaseer):

@Jhannybean - that is not correct

OpenStudy (asnaseer):

@GirlByte - please answer the question I asked above - it will help you work out the answer to this question

OpenStudy (anonymous):

is the side length 24.5 or something

OpenStudy (jhannybean):

\[A = x^2 \iff 49=x^2 \implies x=\sqrt{49}= ~?\]

OpenStudy (asnaseer):

@GirlByte - if a square has a side length of x cm then its area would be \(x^2\) cm2.

OpenStudy (asnaseer):

so you need to find the square root of 49 in order to get the side length

OpenStudy (unanimoose):



OpenStudy (anonymous):

sooo the side length is 7 or am i wrong

OpenStudy (asnaseer):

perfect!

OpenStudy (unanimoose):

Say the side length was j. \[j = \sqrt{49} --> \sqrt{49} \times 2\]

OpenStudy (unanimoose):

Yup.

OpenStudy (asnaseer):

@GirlByte - now put two of these squares together and see if you can work out the perimeter of the resulting rectangle

OpenStudy (jhannybean):